OpenStudy (ltrout):

What is an equation of the line, in point-slope form, that passes through the give point and has the given slope? Points: (8,-8). Slope: (3). Could you explain how to answer it aswell?

OpenStudy (ltrout):

Anyone..?

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (ltrout):

What is the difference between \[y\] And\[y _{1}\] ?

pooja195 (pooja195):

y is just the variable and is part of the formula y1 is the coordinate

OpenStudy (ltrout):

Makes sense now! Thank you :)